# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v9

@@ 4.1 Theorem of Model Interpretability
On the other hand, for any $1\leq i\leq n-1$, suppose there exist some information objects $Iu(t_{i})=\langle ou(t_{i}),Tu_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\rangle$, $Iq(t_{i})=\langle oq(t_{i}),Tq_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\rangle$, and $Ir(t_{i+1})=\langle or(t_{i+1}),Tr_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$, such that the ontological states $S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))$, $S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))$, and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))$ are interpretations in $\mathcal{L}$, and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq\text{Cn}(S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$. Then, according to Theorem 1, $E_{M}=S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))$, $E_{M_{I}}=S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))$, and $E_{M_{O}}=S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))$ are interpretations of $\mathcal{L}$ over the three finite domains $ou(t_{i})\times Tu_{h}(t_{i})$, $oq(t_{i})\times Tq_{h}(t_{i})$, and $or(t_{i+1})\times Tr_{h}(t_{i+1})$, respectively, and $E_{M_{O}}\subseteq\text{Cn}(E_{M}\cup E_{M_{I}})$. At the same time, $S_{M}(M,t_{i})$, $S_{M_{I}}(M_{I},t_{i})$, and $S_{M_{O}}(M_{O},t_{i+1})$ are the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively. By Definition 9, the output $S_{M_{O}}(M_{O},t_{i+1})$ of model M under the input $S_{M_{I}}(M_{I},t_{i})$ is interpretable in the formal system $\mathcal{L}$. Therefore, the sufficiency of the first part of the theorem is established.
@@
For all $i=1,\ldots,n-1$ , by combining Definition 9 with the first part of the theorem, the second part of the theorem follows directly. This completes the proof.
@@ 4.2 Theorem of Ethical Safety Assurance
Ethical safety has long been one of the most widely discussed and scrutinized issues in the field of artificial intelligence mathsea2002 ; Doshi2017 ; Rubin2019 ; Miller2019 , and to date, there is still no universally accepted and practically actionable rigorous definition Russell2015 ; Jobin2019 ; Corrêa2023 . Based on the MLT-MF framework, we are able to theoretically ensure the ethical safety of artificial intelligence through the definition of ethical safety and the corresponding formal expressions. To achieve this, it is first necessary to specify appropriate well-formed formulas $Ec$ as ethical constraints—such as those governing fairness and privacy protection—thus establishing a formalized definition of ethical safety.
Definition 10 (Ethical Safety): Let $S(X,T)$ be the interpretation of a set of well-formed formulas in a formal system $\mathcal{L}$ over the domain $X\times T$. Let $Ec(\vec{v})$ denote an ethical constraint expressed as a well-formed formula in $\mathcal{L}$, where $\vec{v}=(v_{1},\ldots,v_{n})$ are its free variables. For each substitutable $(x,t)\in X\times T$, define
$Ec|_{(x,t)}=Ec(\vec{v}\mapsto\vec{\sigma}(x,t))$ (4.2)
where $\vec{\sigma}:X\times T\to D^{n}$ is a value vector function such that $\sigma_{i}(x,t)\in D$ is the concrete value assigned to $v_{i}$ for $1\leq i\leq n$.
If for all $\varphi\in\text{Cn}(S(X,T))$, and all $(x,t)\in X\times T$,
@@
In the definition of ethical safety, we require not only that $S(X,T)$ contains no formulas that entail $\sim Ec|_{(x,t)}$, but also that its logical closure $\text{Cn}(S(X,T))$ contains no such formulas. This is because if $\text{Cn}(S(X,T))$ were to include formulas entailing $\sim Ec|_{(x,t)}$, then the conjunction of multiple formulas in $S(X,T)$ could still violate the ethical constraint $Ec$. This would indicate that $S(X,T)$ implicitly allows situations that contravene $Ec$, which must be strictly avoided in a rigorous sense. Moreover, it is generally not difficult for a model M to achieve ethical safety with respect to $Ec$ in a simple manner; the key challenge lies in ensuring the largest possible set of valid outputs while maintaining ethical safety. To this end, we propose:
Theorem 5 (Ethical Safety Assurance): Suppose model M is capable of processing $Iq$ with $Iu$ and outputting $Ir$. Let $S_{or}(or,Tr_{h})$ denote the ontological state of $Ir$. For any ethical constraint $Ec$ represented by a well-formed formula in $\mathcal{L}$, there exists a constructive algorithm for obtaining the largest output state subset $S_{or}^{EC}(or,Tr_{h})\subseteq S_{or}(or,Tr_{h})$ that is ethically safe with respect to $Ec$. By introducing an ethical safeguard formula for $Ec$ into $Iu$, it is thereby possible to ensure that model M attains the maximal output state that is ethically safe with respect to $Ec$.
Proof: Since model M is capable of processing $Iq$ with $Iu$ and outputting $Ir$, for the ontological states $S_{ou}(ou,Tu_{h})$, $S_{oq}(oq,Tq_{h})$ and $S_{or}(or,Tr_{h})$ of $Iu$, IqIq and $Ir$ respectively, the MLT-MF framework provides:
$\varphi u=($ $\forall\Phi u)(\forall\Phi q)(\text{IsSubsetOf}^{2}(\Phi u,Sou(ou,Tu_{h}))\land\text{IsSubsetOf}^{2}(\Phi q,S_{oq}(oq,Tq_{h}))\land$
$\text{Processable}^{2}(\Phi u,\Phi q))\to\exists\varphi r(\text{IsFormulaOf}^{2}(\varphi r,S_{or}(or,Tr_{h}))\land$
$\text{Eq}^{2}(\varphi r,f_{process}^{2}(\Phi u,\Phi q)))$ (4.4)
@@
$S_{or}(or,Tr_{h})=\{\varphi r|\varphi r=\text{Processable}^{2}(\Phi u,\Phi q),\Phi u\in Sou(ou,Tu_{h}),\Phi q\in S_{oq}(oq,Tq_{h})\}$ (4.5)
@@
Since $S_{or}(or,Tr_{h})$ is a finite set, we construct a finite hypergraph
$H_{or}=(S_{or}(or,Tr_{h}),\mathcal{E})$ (4.6)
where the vertex set is precisely $S_{or}(or,Tr_{h})$, and the hyperedge set $\mathcal{E}$ covers all well-formed formulas that directly or implicitly violate the ethical constraint $Ec$:
(1) For any $\varphi r\in S_{or}(or,Tr_{h})$ and $(x,t)\in or\times Tr_{h}$, if $\varphi r\rightarrow\sim Ec|_{(x,t)}$, then add the hyperedge $\{\varphi r\}$ to $\mathcal{E}$;
(2) If there exists a minimal subset $W\subseteq S_{or}(or,Tr_{h})$ such that the conjunction of all formulas in $W$ entails $\sim Ec|_{(x,t)}$, then add the hyperedge $W$ to $\mathcal{E}$.
Let $G$ be an independent set in $H_{or}$, i.e., a subset of vertices containing no hyperedge $e\in\mathcal{E}$. If there exists $\varphi\in\text{Cn}(G)$ such that
$\varphi\to\sim Ec|_{(x,t)}$ (4.7)
then there must exist a finite subset $\{\varphi r_{1},\cdots,\varphi r_{k}\}\subseteq G$ such that $\varphi r_{1}\land\cdots\land\varphi r_{k}=\varphi$, and hence $\varphi r_{1}\land\cdots\land\varphi r_{k}\to\sim Ec|_{(x,t)}$. Let $W=\{\varphi r_{1},\cdots,\varphi r_{k}\}$, so the conjunction of $W$ entails $\sim Ec|_{(x,t)}$.
According to the hyperedge construction rules, if $|W|=1$, then $W$ itself directly violates $Ec$, and there must be a hyperedge $\{\varphi r_{1}\}$. But since $\{\varphi r_{1}\}\in G$, this contradicts the assumption that $G$ is an independent set. If $|W|\geq 2$, then by the compactness theorem, the logical entailment can be captured by a finite minimal subset, so there exists a minimal subset $W_{min}\subseteq W$ whose conjunction entails $\sim Ec|_{(x,t)}$. Therefore, there exists a hyperedge $W_{min}\subseteq W\subseteq G$ in $H_{or}$, again contradicting the independence of $G$. Thus, (4.7) cannot hold, and the independent set $G$ is ethically safe with respect to $Ec$.
By the maximum independent set theorem for finite hypergraphs, $H_{or}$ necessarily contains an independent set $G_{max}$ of maximal cardinality. Suppose there exists another ethically safe subset $K$ with
$|K|>|G_{max}|$ (4.8)
then $K$ cannot contain any hyperedge $e\in{E}$. Otherwise, if $|e|=1$, then $K$ directly violates $Ec$; if $|e|\geq 2$, then the formulas in $K$ violate $Ec$. Both cases contradict the ethical safety assumption for $K$. Therefore, $K$ must alsobe an independent set in $H_{or}$, which contradicts the maximality of $G_{max}$. Hence, $G_{max}$ is the largest ethically safe subset of $S_{or}(or,Tr_{h})$ with respect to $Ec$.
@@
In this way, add to $Sou(ou,Tu_{h})$ an ethical safeguard formula:
$\varphi u=($ $\forall\varphi r)(IsFormulaOf^{2}(\varphi r,S_{or}(or,Tr_{h})))\to\exists\varphi s((IsFormulaOf^{2}(\varphi s,S_{os}(os,Ts_{h}))\land$
$(IsFormulaOf^{2}(\varphi r,S_{or}^{Ec}(or,Tr_{h}))\to Eq^{2}(\varphi s,\varphi r))\land$
$(IsFormulaOf^{2}(\varphi r,S_{or}^{\sim Ec}(or,Tr_{h}))\to Eq^{2}(\varphi s,IsSafeback^{1}(NEc))))$ (4.9)
Since the newly added $\varphi u$ pertains to $S_{os}(os,Ts_{h})$ and is entirely independent of the existing formulas in $Sou(ou,Tu_{h})$, it does not affect the satisfaction of Postulate 5 by $Sou(ou,Tu_{h})$. Moreover,
$S_{os}(os,Ts_{h})=S_{or}^{Ec}(or,Tr_{h})\cup\{IsSafeback^{1}(NEc)\}$ (4.10)
Because NEc and IsSafeback1(NEc) were not previously present in $\mathcal{L}$, and $S_{or}^{Ec}(or,Tr_{h})$ and $\{IsSafeback^{1}(NEc)\}$ are logically independent and both logically consistent, $S_{os}(os,Ts_{h})$ is also logically consistent and satisfies all other conditions of Postulate 5. Therefore, it is the maximal output state that is ethically safe with respect to $Ec$.
By taking $S_{os}(os,Ts_{h})$ as the ontological state and mapping it via recoverable information $Is$ to the carrier state $S_{M}(M,Ts_{m})$, $S_{M}(M,Ts_{m})$ is necessarily also ethically safe with respect to $Ec$. This satisfies the condition in Definition 10 for model M to be ethically safe with respect to $Ec$. The theorem is thus proved.
Theorem 5 not only establishes that it is possible to construct an ethical safety assurance mechanism for models, but also provides a constructive algorithm—based on hypergraphs-for obtaining the maximum ethically safe output. However, computing the maximum independent set in a hypergraph is generally an NP-hard problem. In practical applications, the algorithm can be simplified by leveraging two functions: one for ranking the conflict intensity among formulas in $S_{or}(or,Tr_{h})$, and another for ethical safety checking, thus yielding an approximately maximal safe output. This is a practical issue that warrants further investigation and will not be elaborated upon here. Meanwhile, the inheritance of ethical safety in the learning process can be achieved by incorporating dedicated ethical constraint regularization terms into the model algorithm. This involves more specific model structures and algorithmic design, which are beyond the scope of this paper.
@@
4.3 Theorem on Model Generalization Error Estimation
Improving model generalization ability is among the most critical objectives in machine learning. However, in practice, generalization error is usually difficult to analyze and derive theoretically in advance Vapnik1999 ; Bartlett2002 ; Zhang2016 ; Advani2020 ; Perlaza2024 . It can typically only be estimated through testing after training, and it is challenging to obtain accurate error metrics from limited test samples.Applying the MLT-MF framework allows us to estimate upper bounds on the generalization error measured by Total Variation Distance (TVD) between model distributions, facilitating the analysis and prediction of model generalization performance across a variety of tasks.
Theorem 6 (Upper Bound on TVD) Let the information describing the model M after learning and optimization be $I_{u}=\langle ou,Tu_{h},S_{ou},M,Tu_{m},S_{M}\rangle$ and the input information be $I_{q}=\langle oq,Tq_{h},S_{oq},eq,Tq_{m},S_{cq}\rangle$. Assume both $S_{ou}$ and $S_{oq}$ are distributed according to probability distributions over the same measurable space. Let $|S_{oq}|$ and $|S_{ou}\cap S_{oq}|$ denote the cardinalities of the finite sets $S_{oq}$ and $S_{ou}\cap S_{oq}$, respectively. The probability distribution of the model is $p_{M}(x)$, where $p_{M}(x)>0$ for $x\in S_{ou}$, the distribution over $S_{oq}$ is $p_{q}(x)$,with $p_{q}(x)>0$ for $x\in S_{oq}$ . The total probability mass of $p_{q}(x)$ over $S_{ou}\cap S_{oq}$ is denoted $P_{Mq}$. When the model M is sufficiently trained on $S_{ou}$ and adopts a conservative uniform prediction for unknown states, the total variation distance between $p_{M}(x)$ and $p_{q}(x)$ satisfies:
$TVD(p_{q},p_{M})\begin{cases}\approx 0,&S_{oq}\subseteq S_{ou}\\ \leq\sqrt{\frac{1}{2}((1-P_{Mq})\log(\frac{|S_{oq}|-|S_{ou}\cap S_{oq}|}{1-P_{Mq}})-H(p_{q}))},&S_{oq}\neq S_{ou}\cap S_{oq}\end{cases}$ (4.11)
For the second case, it is assumed that $S_{oq}\setminus S_{ou}\neq\emptyset$, $1-P_{Mq}>0$, and $|S_{oq}|-|S_{ou}\cap S_{oq}|>0$. Here, $H(p_{q})=-\sum_{x\in S_{oq}\setminus S_{ou}}p_{q}(x)\log p_{q}(x)$ is the entropy of the data in the non-overlapping region.

### v10

@@ 5.1 Theorem of Model Interpretability
Then, according to Lemma A.1, $E_{M}=S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))$, $E_{M_{I}}=S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i}))$, and $E_{M_{O}}=S_{or}\left(or\left(t_{i+1}\right),{Tr}_{h}\left(t_{i+1}\right)\right)$ are interpretations of $\mathcal{L}$ over the three finite domains $ou(t_{i})\times{Tu}_{h}(t_{i})$, $oq(t_{i})\times{Tq}_{h}(t_{i})$, and $r(t_{i+1})\times{Tr}_{h}(t_{i+1})$, respectively, and $E_{M_{O}}\subseteq Cn(E_{M}\cup E_{M_{I}})$. At the same time, $S_{M}\left(M,t_{i}\right)$, $S_{M_{I}}\left(M_{I},t_{i}\right)$, and $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ are the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively. By Definition 4, the output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ of model $M$ under the input $S_{M_{I}}\left(M_{I},t_{i}\right)$ is interpretable in the formal system $\mathcal{L}$. Therefore, the sufficiency of the first part of the theorem is established.
@@
For all $i=1,\cdots,n-1$, by combining Definition 4 with the first part of the theorem, the second part of the theorem follows directly. This completes the proof.
Definition 4 provides a rigorous mathematical definition of model interpretability. Furthermore, Theorem 1 demonstrates that, throughout the processes of model construction and application, as long as the ontological states of the model and its input and output information can be established within the same formal system, it is possible to reveal the causal relationships between model inputs and outputs within this system, thereby achieving model interpretability. It should also be noted that formal systems vary greatly across different academic disciplines, and the myriad entities in the world can be expressed through information in the forms of formulas, code, text, graphics, or even video. Moreover, the same state can be represented in different information modalities according to varying user requirements. Therefore, Definition 4 and Theorem 1 theoretically establish a foundation for the flexible use of multiple information modalities and for providing model interpretability tailored to user needs.
@@ 5.2 Theorem of Ethical Safety Assurance
To apply the MLT-MF framework, it is first necessary to identify the corresponding well-formed ethical constraint formulas $Ec$ in accordance with fundamental ethical standards such as fairness and privacy protection, thereby establishing a formalized expression of ethical security.
Definition 5 (Ethical Safety) Let $S(X,T)$ be the interpretation of a set of well-formed formulas in a formal system $\mathcal{L}$ over the domain $X\times T$. Let $Ec(\overrightarrow{v})$ denote an ethical constraint expressed as a well-formed formula in $\mathcal{L}$, where $\overrightarrow{v}=(v_{1},\ldots,v_{n})$ are its free variables. For each substitutable $(x,t)\in X\times T$, define
$Ec|_{(x,t)}=Ec(\overrightarrow{v}\mapsto\overrightarrow{\sigma}(x,t))$ (5.2)
where $\overrightarrow{\sigma}:X\times T\rightarrow D^{n}$ is a value vector function such that $\sigma_{i}(x,t)\in D$ is the concrete value assigned to $v_{i}$ for $1\leq i\leq n$.
If for all $\varphi\in Cn(S(X,T))$, and all $(x,t)\in X\times T$,
@@
In the definition of ethical safety, we require not only that $S(X,T)$ contains no formulas that entail $\sim Ec|_{(x,t)}$, but also that its logical closure $Cn(S(X,T))$ contains no such formulas. This is because if $Cn(S(X,T))$ were to include formulas entailing $\sim Ec|_{(x,t)}$, then the conjunction of multiple formulas in $S(X,T)$ could still violate the ethical constraint $Ec$. This would indicate that $S(X,T)$ implicitly allows situations that contravene $Ec$, which must be strictly avoided in a rigorous sense. Moreover, it is generally not difficult for a model M to achieve ethical safety with respect to $Ec$ in a simple manner; the key challenge lies in ensuring the largest possible set of valid outputs while maintaining ethical safety. To this end, we propose:
Theorem 2 (Ethical Safety Assurance) Suppose model M is capable of processing $Iq$ with $Iu$ and outputting $Ir$. Let $S_{or}\left(or,{Tr}_{h}\right)$ denote the ontological state of $Ir$. For any ethical constraint $Ec$ represented by a well-formed formula in $\mathcal{L}$, there exists a constructive algorithm for obtaining the largest output state subset $S_{or}^{Ec}\left(or,{Tr}_{h}\right)\subseteq S_{or}\left(or,{Tr}_{h}\right)$ that is ethically safe with respect to $Ec$. By introducing an ethical safeguard formula for $Ec$ into $Iu$, it is thereby possible to ensure that model M attains the maximal output state that is ethically safe with respect to $Ec$.
Proof: Since model M is capable of processing $Iq$ with $Iu$ and outputting $Ir$, for the ontological states $S_{ou}\left(ou,{Tu}_{h}\right)$, $S_{oq}\left(oq,{Tq}_{h}\right)$ and $S_{or}\left(or,{Tr}_{h}\right)$ of $Iu$, Iq\ Iq and $Ir$ respectively, the MLT-MF framework provides:
$\varphi_{u}\coloneqq$ $\bigl((\forall\,\Phi_{u})(\forall\,\Phi_{q})\bigl({IsSubsetOf}^{2}(\Phi_{u},S_{ou}(ou,{T}_{u,h}))$ (5.4)
$\quad\land{IsSubsetOf}^{2}(\Phi_{q},S_{oq}(oq,{T}_{q,h}))$
$\quad\land{Processable}^{2}(\Phi_{u},\Phi_{q})\bigr)\bigr)$
$\rightarrow\bigl(\exists\,\varphi_{r}\bigl({IsFormulaOf}^{2}(\varphi_{r},S_{or}(or,{T}_{r,h}))$
$\quad\land{Eq}^{2}(\varphi_{r},f_{\mathrm{process}}^{2}(\Phi_{u},\Phi_{q}))\bigr)\bigr).$
@@
$\begin{split}S_{or}\left(or,{Tr}_{h}\right)=\left\{\varphi r\middle|\right.&\left.\varphi r={Processable}^{2}(\Phi u,\Phi q),\right.\\ &\left.\Phi u\in Sou\left(ou,{Tu}_{h}\right),\right.\\ &\left.\Phi q\in S_{oq}\left(oq,{Tq}_{h}\right)\right\}\end{split}$ (5.5)
@@
Since $S_{or}\left(or,{Tr}_{h}\right)$ is a finite set, we construct a finite hypergraph
$H_{or}=\left(S_{or}\left(or,{Tr}_{h}\right),\mathcal{E}\right)$ (5.6)
where the vertex set is precisely $S_{or}\left(or,{Tr}_{h}\right)$, and the hyperedge set $\mathcal{E}$ covers all well-formed formulas that directly or implicitly violate the ethical constraint $Ec$:
(1) For any $\varphi r\in S_{or}\left(or,{Tr}_{h}\right)$ and $(x,t)\in or\times{Tr}_{h}$, if $\varphi r\rightarrow\sim Ec|_{(x,t)}$, then add the hyperedge $\left\{\varphi r\right\}$ to $\mathcal{E}$;
(2) If there exists a minimal subset $W\subseteq S_{or}\left(or,{Tr}_{h}\right)$ such that the conjunction of all formulas in $W$ entails $\sim Ec|_{(x,t)}$, then add the hyperedge $W$ to $\mathcal{E}$.
Let $G$ be an independent set in $H_{or}$, i.e., a subset of vertices containing no hyperedge $e\in\mathcal{E}$. If there exists $\varphi\in Cn(G)$ such that
$\varphi\rightarrow\sim Ec|_{(x,t)}$ (5.7)
then there must exist a finite subset $\left\{{\varphi r}_{1},\cdots,{\varphi r}_{k}\right\}\subseteq G$ such that ${\varphi r}_{1}\land\cdots\land{\varphi r}_{k}=\varphi$, and hence ${\varphi r}_{1}\land\cdots\land{\varphi r}_{k}\rightarrow\sim Ec|_{(x,t)}$. Let $W=\left\{{\varphi r}_{1},\cdots,{\varphi r}_{k}\right\}$, so the conjunction of $W$ entails $\sim Ec|_{(x,t)}$.
According to the hyperedge construction rules, if $|W|=1$, then $W$ itself directly violates $Ec$, and there must be a hyperedge $\left\{{\varphi r}_{1}\right\}$. But since $\left\{{\varphi r}_{1}\right\}\in G$, this contradicts the assumption that $G$ is an independent set. If $|W|\geq 2$, then by the compactness theorem, the logical entailment can be captured by a finite minimal subset, so there exists a minimal subset $W_{\min}\subseteq W$ whose conjunction entails $\sim Ec|_{(x,t)}$. Therefore, there exists a hyperedge $W_{\min}\subseteq W\subseteq G$ in $H_{or}$, again contradicting the independence of $G$. Thus, (4.7) cannot hold, and the independent set $G$ is ethically safe with respect to $Ec$.
By the maximum independent set theorem for finite hypergraphs, $H_{or}$ necessarily contains an independent set $G_{\max}$ of maximal cardinality. Suppose there exists another ethically safe subset $K$ with
$|K|>\left|G_{\max}\right|$ (5.8)
then $K$ cannot contain any hyperedge $e\in\mathcal{E}$. Otherwise, if $|e|=1$, then $K$ directly violates $Ec$; if $|e|\geq 2$, then the formulas in $K$ violate $Ec$. Both cases contradict the ethical safety assumption for $K$. Therefore, $K$ must also be an independent set in $H_{or}$, and (5.8) cannot hold. Hence, $G_{\max}$ is the largest ethically safe subset of $S_{or}\left(or,{Tr}_{h}\right)$ with respect to $Ec$.
@@
In this way, add to $Sou\left(ou,{Tu}_{h}\right)$ an ethical safeguard formula:
$\begin{split}\varphi u=&(\forall\varphi r)({IsFormulaOf}^{2}\left(\varphi r,S_{or}\left(or,{Tr}_{h}\right)\right))\rightarrow\exists\varphi s\left({IsFormulaOf}^{2}\left(\varphi s,S_{os}\left(os,{Ts}_{h}\right)\right)\right.\\ &\left.\land\left({IsFormulaOf}^{2}\left(\varphi r,S_{or}^{Ec}\left(or,{Tr}_{h}\right)\right){\rightarrow Eq}^{2}(\varphi s,\varphi r)\right)\right.\\ &\left.\land\left({IsFormulaOf}^{2}\left(\varphi r,S_{or}^{\sim Ec}\left(or,{Tr}_{h}\right)\right){\rightarrow Eq}^{2}\left(\varphi s,{IsSafeback}^{1}(NEc)\right)\right)\right)\end{split}$ (5.9)
Since the newly added $\varphi u$ pertains to $S_{os}\left(os,{Ts}_{h}\right)$ and is entirely independent of the existing formulas in $Sou\left(ou,{Tu}_{h}\right)$, it does not affect the satisfaction of Postulate 5 by $Sou\left(ou,{Tu}_{h}\right)$. Moreover,
$S_{os}\left(os,{Ts}_{h}\right)=S_{or}^{Ec}\left(or,{Tr}_{h}\right)\cup\left\{{IsSafeback}^{1}(NEc)\right\}$ (5.10)
Because NEc and ${IsSafeback}^{1}(NEc)$ were not previously present in $\mathcal{L}$, and $S_{or}^{Ec}\left(or,{Tr}_{h}\right)$ and $\left\{{IsSafeback}^{1}(NEc)\right\}$ are logically independent and both logically consistent, $S_{os}\left(os,{Ts}_{h}\right)$ is also logically consistent and satisfies all other conditions of Postulate A.1. Therefore, it is the maximal output state that is ethically safe with respect to $Ec$.
By taking $S_{os}\left(os,{Ts}_{h}\right)$ as the ontological state and mapping it via recoverable information $Is$ to the carrier state $S_{M}(M,{Ts}_{m})$, $S_{M}(M,{Ts}_{m})$ is necessarily also ethically safe with respect to $Ec$. This satisfies the condition in Definition 5 for model M to be ethically safe with respect to $Ec$. The theorem is thus proved.
Theorem 2 not only establishes that it is possible to construct an ethical safety assurance mechanism for models, but also provides a constructive algorithm—based on hypergraphs—for obtaining the maximum ethically safe output. However, computing the maximum independent set in a hypergraph is generally an NP-hard problem. In practical applications, the algorithm can be simplified by leveraging two functions: one for ranking the conflict intensity among formulas in $S_{or}\left(or,{Tr}_{h}\right)$, and another for ethical safety checking, thus yielding an approximately maximal safe output. This is a practical issue that warrants further investigation and will not be elaborated upon here. Meanwhile, the inheritance of ethical safety in the learning process can be achieved by incorporating dedicated ethical constraint regularization terms into the model algorithm. This involves more specific model structures and algorithmic design, which are beyond the scope of this paper.
@@
5.3 Theorem on TVD Upper Bounds
Applying the MLT-MF framework allows us to estimate upper bounds on the generalization error measured by Total Variation Distance (TVD) between model distributions, facilitating the analysis and prediction of model generalization performance across a variety of tasks.
Theorem 3 (TVD Upper Bound for Conservative Prediction Strategies in Noise-Free Inputs and Uncovered Training Regions) Let the information describing the model $M$ after learning and optimization be $Iu=\left\langle ou,{Tu}_{h},S_{ou},M,{Tu}_{m},S_{M}\right\rangle$, and the input information be $Iq=\left\langle oq,{Tq}_{h},S_{oq},cq,{Tq}_{m},S_{cq}\right\rangle$. Assume both $S_{ou}$ and $S_{oq}$ are distributed according to probability distributions over the same measurable space. Let $\left|S_{oq}\right|$ and $\left|S_{ou}\cap S_{oq}\right|$ denote the cardinalities of the finite sets $S_{oq}$ and $S_{ou}\cap S_{oq}$, respectively. The probability distribution of the model is $p_{M}(x)$, where $p_{M}(x)>0$ for $x\in S_{ou}$ ; the distribution over $S_{oq}$ is $p_{q}(x)$, with $p_{q}(x)>0$ for $x\in S_{oq}$. The total probability mass of $p_{q}(x)$ over $S_{ou}\cap S_{oq}$ is denoted $P_{Mq}$. When the model $M$ is sufficiently trained on $S_{ou}$, and adopts a conservative uniform prediction for unknown states, the total variation distance between $p_{M}(x)$ and $p_{q}(x)$ satisfies:
$TVD\left(p_{q},p_{M}\right)$ $=\begin{cases}\approx 0,&\text{if }S_{oq}\subset S_{ou}\\[5.0pt] \leq\sqrt{\frac{1}{2}((1-P_{Mq})\log{((\left|S_{oq}\right|-\left|S_{ou}\cap S_{oq}\right|)}/(1-P_{Mq}))-H(p_{q}))},&\text{if }S_{oq}\neq S_{ou}\cap S_{oq}\end{cases}$ (5.11)